OpenStudy (anonymous):

A line passes through the point (0, –4) and has a slope 3/2 of What is the equation of the line? A.3x-2y=-8 B.6x+4y=-8 C. 2x+4y=-1 D.3x-2y=8

OpenStudy (ddcamp):

First, solve for point-slope form: \[y-y_1=m(x-x_1)\] Where m is the slope and (x1,y1) is the point it passes through.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{3}{2} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

well, or don't solve by "y" just set it in "standard form"

OpenStudy (anonymous):

Your numbers confuse me